OpenStudy (anonymous):

Solve the system of the quadratic equations : y=x^2-2x+2 Y-2x=-2

OpenStudy (anonymous):

You can solve by substitution: \[y-2x=-2~~\Rightarrow~~y=2x-2\] Substituting into the quadratic, you have \[2x-2=x^2-2x+2\\ x^2-4x+4=0\\ (x-2)^2=0\] Solve for \(x\), then \(y\).

OpenStudy (anonymous):

How do I solve for x and y @sithsandgiggles

OpenStudy (anonymous):

From the last line, \((x-2)^2=0\), this is only true for \(x=2\). Plug this into either original equation to solve for \(y\). The nonquadratic equation is simpler: \[y-2x=-2~~\Rightarrow~~y-2(2)=-2~~\Rightarrow~~y=\cdots\]

OpenStudy (anonymous):

Y=2? @sithsandgiggles

OpenStudy (anonymous):

Yep